Use the Substitution method to solve the system of equations.
3x + y = 5
4x - 7y = -10

Answer by checkley77(12844)

3x+y=5
Y=-3X+5
4x-7y=-10 REPLACE Y IN THIS EQUATION BY (-3X+5)
4X-7(-3X+5)=-10
4X+21X-35=-10
25X=-10+35
25X=25
X=25/25=1 ANS.
3*1+Y=5
3+Y=5
Y=5-3
Y=2 ANS.
PROOF:
4*1-7*2=-10
4-14=-10
-10=-10